OpenStudy (anonymous):

m13 × m6 = m19 m7 m18 m8

Mehek (mehek14):

is this supposed to be m to the power 13 and m to the power 6?

OpenStudy (anonymous):

m19. Because you are only combining like terms right now.

Mehek (mehek14):

if it is, then when you multiply, you add the exponents

OpenStudy (anonymous):

yess m to the power of (13) and (m) to the power of 6

OpenStudy (mathstudent55):

I assume those are exponents, right? \(m^{13} \times m^6\) Here is the rule: \(a^m \times a^n = a^{m + n} \) When you multiply together powers with the same base, you add the exponents.

OpenStudy (anonymous):

thank you @mathstudent55